A 6-pole generator has a lap-wound armature with 40 slots with 20 conductors per slot.The flux per pole is 25mWb.The terminal voltage is 288v and the voltage regulation of the generator is 4%. Determine the speed of the generator.


1
Expert's answer
2021-05-07T08:12:55-0400

P=6

A=P

Z=20*40=800

Flux per pole = 25mWb

V=288v

voltage regulation of the generator = 4%=0.04

voltage regulation of the generator=(V-E)/V

E=276.48 volt

E=(N*P*Flux per pole*Z)/A

N=13.824 rps

N=829.44 rpm